OpenStudy (anonymous):

how to solve this quadratic equation by completing perfect square a^2.u^2+2abcd.u-(1+2c)b^2.d^2=0

OpenStudy (anonymous):

\[a^2u^2+2abcd.u-(1+2c)b^2d^2=0\]

OpenStudy (mathstudent55):

The variable is u, and all other letters are constants?

OpenStudy (anonymous):

i did it till \[(u+\frac{ bcd }{ a })^2=\frac{ (1+2c)b^2d^2 }{ a^2 }+(\frac{ bcd }{ a })^2\]

OpenStudy (mathstudent55):

Let me try. To compete the square, the leading coefficient must be 1, so if it isn't 1, we divide the entire equation by the leading coefficient.

OpenStudy (mathstudent55):

\(\large a^2u^2+2abcdu-(1+2c)b^2d^2=0\) \(\large u^2+\dfrac{2abcdu}{a^2}-\dfrac{(1+2c)b^2d^2}{a^2}=0 \) \(\large u^2+\dfrac{2bcdu}{a}-\dfrac{(1+2c)b^2d^2}{a^2}=0 \)

OpenStudy (mathstudent55):

Now we move the constant term to the right side: \(\large u^2+\dfrac{2bcdu}{a} = \dfrac{(1+2c)b^2d^2}{a^2}\)

OpenStudy (mathstudent55):

Now we add the square of half of the second term's coefficient to both sides. \(\large u^2+\dfrac{2bcd}{a}u + \left(\dfrac{bcd}{a}\right)^2 = \dfrac{(1+2c)b^2d^2}{a^2} + \left(\dfrac{bcd}{a}\right)^2 \)

OpenStudy (anonymous):

i have written the next step

OpenStudy (mathstudent55):

\(\large \left(u +\dfrac{bcd}{a}\right)^2 = \dfrac{(1+2c)b^2d^2 +b^2c^2d^2}{a^2} \)

OpenStudy (mathstudent55):

So far we have the same.

OpenStudy (mathstudent55):

Now take roots and don't forget +/- on the right side. I am assuming the constants are nonnegative. \(\large u +\dfrac{bcd}{a} = \pm \sqrt{\dfrac{(1+2c)b^2d^2 +b^2c^2d^2}{a^2}}\) \(\large u +\dfrac{bcd}{a} = \pm \dfrac{bd}{a}\sqrt{(1+2c) +c^2}\)

OpenStudy (anonymous):

\[\sqrt{(1+2c)+c^2}=\sqrt{(c+1)^2}\]

OpenStudy (anonymous):

so \[u+\frac{ bcd }{ a }=\pm[\frac{ bcd }{ a }+\frac{ bd }{ a }]\]

OpenStudy (anonymous):

so i think the answer is \[u=\frac{ bd }{ a }\] and \[u=\frac{ -bd(2c+1) }{ a }\]

OpenStudy (mathstudent55):

\(\large u =-\dfrac{bcd}{a} \pm \dfrac{bd}{a}\sqrt{(1+2c) +c^2}\) \(\large u =\dfrac{-bcd \pm bd\sqrt{c^2 + 2c + 1}}{a}\) \(\large u =\dfrac{-bdc \pm bd\sqrt{(c + 1)^2}}{a}\) \(\large u =\dfrac{-bdc \pm bd(c + 1)}{a}\) \(\large u =\dfrac{-bdc \pm (bdc + bd)}{a}\) \(\large u =\dfrac{bd}{a}\) or \(\large u =\dfrac{-2bdc - bd}{a}\) \(\large u =\dfrac{bd}{a}\) or \(\large u =\dfrac{-bd(2c + 1)}{a}\)

OpenStudy (mathstudent55):

Yes, you are correct.

OpenStudy (anonymous):

THANK YOU

OpenStudy (mathstudent55):

You're welcome.